the The net area of the region between y = x3 - 2x and the over curve x-axis interval [0, 2] is
Answer #1

Y=x² zx X-2 x²2x=0 x(x²2)=0 x=0,22 52 Area=- [622-2x)dx + [@2x)dx - x² 0 4 52 + o 1272 S2 2 472 - 2 + 2 = (2-1)+(A_4-1 +2) =1